In relation to producers, consumers and competitive markets, the domestic supply and demand curves for pecan nuts in South Africa at equilibrium are Supply P=50+Q
Demand P=200-2Q
Where P is the price in rand per kilogram and Q is the quantity in millions of kilograms
1
Expert's answer
2020-11-10T10:34:18-0500

In equilibrium Qd = Qs, so:

200 - 2Q = 50 + Q,

3Q = 150,

Q = 50 millions of kilograms,

P = 50 + 50 = 100 rands.